### v1

# SySLLM: Generating Synthesized Policy Summaries for Reinforcement Learning Agents Using Large Language Models

Abstract
Policies generated by Reinforcement Learning (RL) algorithms can be difficult to describe to users, as they result from the interplay between complex reward structures and neural network-based representations. This combination often leads to unpredictable behaviors, making policies challenging to analyze and posing significant obstacles to fostering human trust in real-world applications. Global policy summarization methods aim to describe agent behavior through a demonstration of actions in a subset of world-states. However, users can only watch a limited number of demonstrations, restricting their understanding of policies. Moreover, those methods overly rely on user interpretation, as they do not synthesize observations into coherent patterns. In this work, we present SySLLM (Synthesized Summary using LLMs), a novel method that employs synthesis summarization, utilizing large language models’ (LLMs) extensive world knowledge and ability to capture patterns, to generate textual summaries of policies. Specifically, an expert evaluation demonstrates that the proposed approach generates summaries that capture the main insights generated by experts while not resulting in significant hallucinations. Additionally, a user study shows that SySLLM summaries are preferred over demonstration-based policy summaries and match or surpass their performance in objective agent identification tasks.  [^1] All relevant resources for this study, such as SySLLM summaries, HIGHLIGHTS summaries, expert summaries, experience videos, the source code, and other materials are available at https://osf.io/c78jb/?view_only=76fbf59131d4409bac222e5929a654f2.
1 Introduction
Explainable Reinforcement Learning (XRL) methods aim to enable human users to understand agents’ policies and decision-making processes. Such understanding is crucial for fostering trust and ensuring the real-world deployment of RL agents. Several approaches to XRL have been developed in recent years [25], including saliency maps that visualize the agent’s attention, reward bars that break down expected utility into components, and causal explanations that clarify why actions were taken. Other methods extract logical rules to describe when actions are taken, but these approaches rely on predefined logical representations and often focus on individual actions rather than broader behavior patterns. Despite these advancements, experiments reveal that human users still struggle to form accurate mental models of agent behavior [6, 19, 23].
Recent advancements in large, pre-trained language models (LLMs) present a novel opportunity to enhance explanation methods for reinforcement learning (RL) policies. These models exhibit exceptional in-context learning abilities and possess extensive commonsense knowledge about human behavior, which makes them promising tools for generating meaningful, language-based explanations of RL policies. LLMs have already been employed in constructing world models and planning with domain-independent planners, showcasing their potential for reasoning and providing explanation [14].
Figure 1: Collecting the experience dataset (Section 4.2).
In this paper, we aim to leverage LLMs to generate comprehensive global explanations of RL agent policies. Specifically, our goal is to use LLMs to produce natural language summaries that encapsulate key patterns of behavior, capabilities, and limitations of RL policies. This approach seeks to build upon previous work that utilized demonstration-based policy summaries [4], where agent policies are described through selected trajectories in video. However, these methods are inherently constrained by the representativeness of the selected scenarios and the limited number of demonstrations that can feasibly be observed. They do not synthesize observations into coherent summaries, leaving users to rely heavily on their own interpretations to discern patterns. In contrast, we hypothesize that LLM-generated summaries can effectively distill these patterns into concise, coherent narratives, providing users with a clearer and more accessible understanding of agent policies.
Leveraging LLMs to generate accurate and comprehensible policy summaries is a non-trivial task. While LLMs exhibit remarkable reasoning abilities, they are prone to hallucinations—producing outputs that seem plausible but deviate from the agent’s true behavior [16]. This raises questions about whether such descriptions can reliably reflect the nuances of RL agent policies. Moreover, their utility heavily depends on end users’ ability to interpret them. Laypeople, who often lack familiarity with RL concepts, may struggle to form accurate mental models of agent behavior. Natural language explanations, while accessible, risk oversimplifying or failing to convey subtle patterns [6].
A further challenge lies in capturing the spatial and temporal relationships inherent in RL environments. These relationships evolve continuously over many steps and episodes: spatial dependencies include changing object positions, while temporal dependencies arise from sequential actions. Since LLMs are trained primarily on static text, they may struggle to model these dynamics effectively [24]. Together, these factors complicate the task of distilling and communicating agent behaviors in a way that remains faithful to the policy while being accessible to users.
Our method, Synthesized Summary using LLMs (SySLLM), leverages LLMs to generate comprehensive summaries of RL agent behavior. The proposed pipeline first collects the agent’s interactions within the environment (Figure 1) and converts these into natural language descriptions of observations and actions. These descriptions are then used to prompt an LLM, which synthesizes a natural language summary that captures key patterns in the agent’s behavior (Figure 2). We implement this methodology across five distinct environments, generating language-based summaries for agents with diverse behavioral patterns.
To evaluate SySLLM, we employ two complementary methods. First, we compare its summaries to those created by six RL experts, finding substantial agreement between the two. Second, we conduct a user study where participants interact with two types of policy summaries: visual summaries generated by the HIGHLIGHTS-DIV algorithm [3] and language-based summaries generated by SySLLM. Participants rated SySLLM higher in subjective preferences and performed equally well or better at identifying agents based on these summaries. These results highlight the effectiveness of SySLLM in both comprehensibility and practical utility.
The main contributions of our work are:
  • We introduce Synthesized Summary using LLMs (SySLLM), a novel and versatile method that generates holistic, synthesized, language-based summaries of RL agent policies using LLMs.
  • We validate the correctness of SySLLM by comparing its summaries to those of RL experts, showing substantial alignment and demonstrating that SySLLM accurately captures key policy details.
  • We conduct a large-scale user study, demonstrating that participants prefer SySLLM over a state-of-the-art demonstration-based policy summarization method in a summary preference task and perform equally well or better in an agent identification task.
Figure 2: Generating natural language global policy summaries (Section 4.3).
2 Background and Related Work
Explainable RL (XRL)
Our work extends the body of research concerned with explaining RL policies and particularly the line of work on global policy explanations (see Amitai and Amir [5] for a recent survey). Policy summarization approaches present users with demonstrations of the agent acting in the environment to convey its behavior. Several criteria for extracting such summaries have been proposed in the literature, including state importance, diversity, visitation frequency, and policy reconstruction [3, 17, 23, 18, 34]. While these approaches were shown to improve people’s ability to assess agent capabilities and predict agent behavior, users can only grasp a limited amount of demonstrations, and experiments show that their interpretation of policies highly depends on the subset of examples shown to them [23].
Other works propose learning a simplified model of the policy (e.g., a decision tree or a set of logical rules) that captures the essence of the agent’s decision-making [38, 8, 7]. Yet, these approaches are also limited due to the tradeoff between the accuracy of the simplified model and its interpretability to users.
We hypothesize that LLMs may offer a complementary approach to global explanations by capturing meaningful patterns of agent behavior and presenting them in understandable language to users.
Pre-trained Language Models in RL
Large Language Models (LLMs) are probabilistic text models trained on substantial text datasets. Models like GPT [30], which are autoregressive, excel in text generation by sequentially predicting future words based on the preceding text. Such models can be useful in the context of RL as they encode common-sense knowledge about the world. Conversely, models such as BERT [12], which are based on the masked framework, focus on understanding deep sentence semantics by training to recognize parts of text that are intentionally obscured. Encoder-decoder models merge these traits to adeptly handle tasks requiring both the interpretation and creation of text. These models have been extensively pre-trained on large-scale text corpora, equipping them with the ability to tackle complex language processing tasks with minimal specific instruction, often achieving impressive results [9]. This extensive training allows them to efficiently process and summarize complex narratives, and to identify relevant patterns within large datasets.
Recent work has developed methods that use LLMs in the training of RL policies. For example, LLMs are being used to guide intrinsic exploration by suggesting plausibly useful goals and providing intrinsic rewards [13, 27]. Other works use LLMs to generate useful state abstractions [29, 28]. Additionally, LLMs are employed to provide natural language instructions and dynamic feedback, enhancing agent learning in complex, open-ended environments [39]. Differently from these approaches, in this work, we aim to use LLMs to explain RL policies.
3 Problem Statement
We define a partially observed Markov Decision Process (POMDP) as a tuple $\mathcal{M}=\langle\mathcal{S},\mathcal{A},\mathcal{O},T,O,R,\gamma\rangle$, where observations $o\in\mathcal{O}$ are derived from states $s\in\mathcal{S}$ and actions $a\in\mathcal{A}$ via $O(o|s,a)$. The dynamics are governed by the transition function $T(s^{\prime}|s,a)$, with $R$ as the reward function and $\gamma$ as the discount factor. POMDPs generalize Markov Decision Processes (MDPs), where observations fully reveal the state.
The agent operates under a stochastic policy $\pi:\mathcal{O}\rightarrow\Delta(\mathcal{A})$, mapping each observation $o\in\mathcal{O}$ to a probability distribution over actions $\mathcal{A}$. The objective of the policy is to maximize the expected cumulative discounted reward.
Given a POMDP environment and the agent’s policy, our goal is to generate a natural language summary that encapsulates the agent’s strategy and decision-making process. This summary explains the policy’s operational dynamics, effectiveness, and behavior across scenarios. Formally, a summarization is a mapping from $(\mathcal{M},\Pi)$ to $\mathcal{T}$, where $(\mathcal{M},\Pi)$ represents environment-policy pairs, and $\mathcal{T}$ is the set of natural language summaries.
4 Methodology
Our method consists of several key components, which we formalize in Algorithm 1. In the first phase (illustrated in Figure 1), we run the agent in its environment and perform a captioning procedure to convert experiences into natural language. These captioned experiences are logged into an experience dataset (Sections 4.1 and 4.2). In the second phase, which we illustrate in Figure 2, we take the experience dataset as input and output a natural language summary that aims to capture patterns in the agent’s policy (Section 4.3). To do so, we compose a prompt with the formatted experience dataset that is then sent to an LLM to generate the summary.
4.1 Captioner
As will be detailed in Section 4.2, our approach utilizes captioners to leverage LLMs for summarizing an agent’s policy. We convert each observation $o\in\mathcal{O}$ and action $a\in\mathcal{A}$ into natural language descriptions that highlight their key features. In this subsection, we formalize the role and function of the captioners.
Observation Captioner:
The observation captioner, $C_{\text{observation}}$, translates observations into natural language descriptions:
$C_{\text{observation}}:\mathcal{O}\rightarrow\Sigma^{\star}.$
Where $\Sigma^{\star}$ represents the set of all possible strings. This function distills complex state information into comprehensible text, such as translating a grid map into a description of object locations.
Action Captioner:
Similarly, the action captioner, $C_{\text{action}}$, maps actions to natural language representations:
$C_{\text{action}}:\mathcal{A}\rightarrow\Sigma^{\star}.$
Similar to past work [26, 20] we assume the availability of a captioner language annotator. The process of generating captions is supported by a range of established methods, which are further discussed in Section 8.
4.2 Collecting the Agent’s Experience Dataset
We describe the first phase of Algorithm 1, detailed in Lines 114.
We assume that the environment ($ENV$) supports two functions. The first function is RESET, which resets the environment to its initial state and returns the initial observation, and STEP, which takes an action as input and returns the next observation, the reward for that action, a boolean indicating whether the episode has ended, and additional information.
First, we initialize the experience dataset $ED$ (Line 1). The for loop (Line 2) iterates over multiple episodes, initializing the step counter $t=1$ at the beginning of each episode (Line 3) and obtaining the initial observation $o_{1}$ (Line 4). The variables $epReward$ and $done$ are also initialized to track the cumulative reward and episode completion status, respectively (Lines 56). The while loop (Line 7) continues until the episode concludes.
During each iteration of the while loop, an action is sampled from the policy $\pi$ (Line 8), and both the observation and action are converted into natural language representations by the captioners. These representations are added to the experience dataset along with the cumulative reward (Line 9). Subsequently, the environment is updated using the selected action, and the next observation is retrieved (Line 10).
Finally, the cumulative reward and step index are updated accordingly (Lines 1112). The information stored in the experience dataset is detailed in Section B of the Appendix.
4.3 Generating Natural Language Summaries
We move on to addressing the second phase of Algorithm 1, covered in Lines 1517.
To generate natural language summaries, we use the experience dataset and an LLM. The process begins by formatting the experience dataset into a structured format suitable for integration into a prompt. This format captures the sequential and contextual details of the agent’s interactions, creating a comprehensive narrative of its decision-making process (Line 15).
The prompt design, inspired by Chain-of-Thought reasoning [40], decomposes complex behavioral analysis into specific components. By incorporating detailed environmental parameters and quantitative evaluation guidelines, the prompt effectively guides the model’s analysis, enhancing its ability to generate insightful summaries of reinforcement learning behavior. This design was refined through iterative exploration, as detailed in Section H of the Appendix.
The prompt structure consists of four main components:
  • General Instructions: for example, “Generate a summary that captures the essence of the agent’s policy.”
  • Environment Description: such as, “Describe the agent’s task and key aspects of the environment.”
  • Interpretation Instructions: e.g., “Descriptions are from the agent’s perspective.”
  • Output Instructions: for instance, “Provide a concise summary.”
The LLM then produces the summary based on the prompt and the formatted experience dataset (Line 16).
Algorithm 1 SySLLM algorithm
Input: Environment $ENV$, trained policy $\pi$, captioners $C_{\text{observation}}$, $C_{\text{action}}$
Parameter: number of episodes $N$, prompt $P$
Output: summary $\in\mathcal{T}$
1:  Initialize experience dataset $ED$
2:  for $i=1$ to $N$ do
3:     Let $t\leftarrow 1$
4:     $o_{t}\leftarrow$ $ENV$.RESET()
5:     $epReward\leftarrow 0$
6:     $done\leftarrow\text{False}$
7:     while $\neg done$ do
8:        $a_{t}\sim\pi(\cdot\mid o_{t})$
9:        $ED$.ADD($C_{\text{observation}}(o_{t})$, $C_{\text{action}}(a_{t})$, $epReward$)
10:        $o_{t+1},r_{t+1},done\leftarrow$ $ENV$.STEP($a_{t}$)
11:        $epReward\leftarrow epReward+r_{t+1}$
12:        Let $t\leftarrow t+1$
13:     end while
14:  end for
15:  $formattedInput$leftarrow $P$ + FORMAT($ED$)
16:  $summary$leftarrow LLM($formattedInput$)
17:  return $summary$
5 Implementation
We demonstrate our approach in five diverse environments to evaluate summaries for seven agents, each with a unique policy. Specifically, we use the MiniGrid framework [10].
5.1 Captioners
To convert raw observations and actions into natural language, we designed a rule-based captioning system for the agent’s grid-based representation of the environment. This system systematically generates detailed descriptions of visible elements, emphasizing spatial relationships for navigation. Each action in the environment’s discrete action space is translated into natural language annotations, such as ‘move forward’ or ‘turn left’. We evaluated the effectiveness of this captioner by experimenting with and refining its design, ensuring the generated descriptions are accurate and comprehensive.
5.2 Agents’ Policy Training
We trained three qualitatively different agents for the MiniGrid-Unlock environment:
  • Goal-directed Agent: Follows a policy designed to minimize the number of steps to unlock the door efficiently. The agent can see all grid cells within a 180-degree field of vision in front of it.
  • Short-Sighted Agent: Follows a policy designed to minimize the number of steps to unlock the door efficiently but the agent’s sight is $3\times 3$.
  • Random Agent: Selects actions uniformly from the action space. We chose this policy to see how the LLM summarization process handles cases where there is no clear pattern of behavior.
Additionally, we trained four other agents in the following MiniGrid environments: Dynamic Obstacles, Lava Gap, Red-Blue Doors, and Crossing. Using SySLLM, we generated a policy summary for each of these agents. All of the agents are trained using the PPO algorithm [33, 31]. An overview of agent performance is shown in Table 1. The mean reward for the first three agents in the Unlock environment shows a declining trend, which aligns with the policies that each of these agents employ. In addition, the success rate of the agents in the other environments remains consistently high. Implementation details are provided in Section E of the Appendix.
Agent Mean Reward ± SD Mean Length Success Rate
Unlock Goal-directed 0.73 ± 0.21 20.25 0.93
Unlock Short-sighted 0.41 ± 0.27 44.43 0.77
Unlock Random 0.00 ± 0.01 70.00 0.00
Dynamic Obstacles 0.78 ± 0.06 17.20 1.00
Lava Gap 0.896 ± 0.02 10.82 1.00
Red Blue Doors 0.70 ± 0.26 17.06 0.88
Crossing 0.67 ± 0.18 24.80 0.94
Table 1: Report of agents’ performance metrics for 500 episodes.
5.3 Generating Policy Summaries
For each agent, we collect an experience dataset that logs the captioned observations, actions taken by the agent, and additional relevant data. Specifically, we log 50 episodes for each agent. After collecting and formatting the data, we synthesized the narratives into coherent global policy summaries for each agent using zero-shot prompting. We employed the gpt-4-turbo model with a temperature of $0.2$ [1]. The prompt used is shown in Section C of the Appendix.
In selecting the final summary, we employed a methodical approach by first generating ten distinct summaries through Line 16 of our algorithm. Each of these summaries was then embedded into a high-dimensional vector space using text-embedding-3-small, allowing us to quantify the semantic proximity of each summary to the others within this space. The centroid of these embeddings was calculated to identify the geometric mean of the summaries, representing the central tendency of the semantic dimensions explored by the model.
We chose the summary that is the median (5th) closest to the centroid. This position represents a solution to the trade-off between generality and specificity. Summaries closer to the centroid tend to include more general statements that apply broadly across different scenarios, while those further away incorporate more specific details unique to particular contexts. An example of a selected summary is provided in Section D of the Appendix.
5.4 Insights From Summaries
Using the SySLLM method, we generated summaries capturing key aspects of each agent’s decision-making and behavior. This section presents key insights from these summaries, concisely illustrating agents’ policies and critical action patterns. Table 2 provides a overview of these insights, showcasing their alignment with task objectives and how agents behave across various environments.
For the goal-directed agent in the Unlock environment, Insight (2) states that the agent “demonstrates a pattern of turning towards the nearest key or door once identified, suggesting a straightforward heuristic of minimizing distance to the target.” This reflects the agent’s efficient and consistent behavior, dynamically adjusting its path to achieve objectives with minimal steps. Insight (3) quantifies this efficiency, noting that episodes are completed within 15-25 steps on average, often achieving near-maximum rewards. These observations are validated by performance metrics in Table 1, which report a mean episode length of 20.25 and a high mean reward of 0.71. Together, these results highlight SySLLM’s ability to capture both qualitative patterns and quantitative performance metrics.
In the Dynamic Obstacles environment, SySLLM captures the agent’s ability to navigate challenges effectively. Insight (1) states that the agent “consistently demonstrates the ability to avoid obstacles (blue balls) by making turns or moving forward when the path is clear.” This highlights its proficiency in avoiding collisions, a critical success factor. However, Insight (3) notes inconsistent behavior when navigating closely spaced obstacles, with the agent occasionally taking longer routes or making unnecessary turns. These observations demonstrate SySLLM’s ability to provide a nuanced understanding of both the agent’s strengths in straightforward scenarios and its struggles in more complex configurations.
Agent Insights from Summaries
Unlock
Goal-Directed
(1) The agent effectively identifies keys and adjusts its path based on their relative position, shifting focus to unlocking the door.
(2) It consistently turns towards the nearest key or door, minimizing distance, which remains consistent across episodes.
(3) The agent completes episodes efficiently, averaging 15-25 steps with near-maximal cumulative rewards.
Unlock
Short-sighted
(1) The agent follows a right-wall method, moving forward until encountering an obstacle before turning.
(2) It identifies keys and doors efficiently, maneuvering toward and using them correctly.
(3) Decisions are heavily influenced by its immediate field of vision, reacting only to nearby objects.
Unlock Random (1) The agent exhibits unstructured behavior, often repeating unnecessary actions.
(2) It frequently toggles doors multiple times or picks up and drops keys without using them effectively.
Lava Gap (1) The agent consistently avoids lava, demonstrating awareness of environmental hazards.
(2) Upon encountering an obstacle, it either turns or moves in the opposite direction.
Red-Blue Doors (1) The agent prioritizes opening the red door before the blue door, optimizing reward accumulation.
(2) It successfully interacts with doors in a structured sequence, adhering to task constraints.
Crossing (1) The agent moves towards the green goal once it enters its field of vision, adjusting its path accordingly.
(2) It avoids collisions with walls through timely directional changes.
Dynamic
Obstacles
(1) The agent effectively avoids moving obstacles (blue balls) by adjusting its movement.
(2) It identifies objects in its field of vision and makes informed navigation decisions.
(3) In dense obstacle scenarios, occasional inefficiencies or unnecessary turns are observed.
Table 2: Summarized insights on agent behaviors across different tasks.
6 Expert Evaluation of Summaries
To validate the quality and alignment of SySLLM summaries, we conducted evaluations with RL experts. Specifically, six graduate students actively engaged in RL research with experience in training agents were recruited. These experts were divided into two groups: Experts 1, 2, and 3 evaluated the Minigrid-Unlock goal-directed and short-sighted agents, while Experts 4, 5, and 6 evaluated the remaining agents. Each expert was shown videos approximately 120 seconds long, depicting the agents’ behaviors in the environment. Based on these videos, the experts were tasked with summarizing the agents’ behaviors. To ensure a fair comparison with SySLLM-generated summaries, we provided the experts with instructions that aligned closely with the prompts used for the language model (see Section F of the Appendix for detailed instructions).
To evaluate how much of the experts’ observations were captured by the SySLLM summaries, we used a recall metric. This metric quantified the extent to which the SySLLM summary covered the points identified by the experts. Each expert summary and the SySLLM summary were broken down into lists of key points, and a scoring system was applied: 1 point for a match with a SySLLM key point, 0.5 points for a partial match, and 0 points for no match. The recall scores, as shown in Table 3, ranged from 0.687 for the Unlock goal-directed agent, indicating moderate agreement, to 0.914 for the Crossing agent, reflecting a near-complete capture of expert insights. The overall mean recall score across all agents was 0.814, demonstrating substantial coverage of the experts’ points by the SySLLM summaries.
In the second evaluation phase, the experts were presented with key points from the SySLLM summaries that they had not mentioned in their own summaries. They were tasked with classifying these points as Matched, Partially Matched, or Not Matched based on their agreement with the content. Using the same scoring system as for recall (Matched = 1, Partially Matched = 0.5, Not Matched = 0), we calculated a precision score for each summary. The precision score measures how many of the SySLLM summary points were validated as correct by the experts. Importantly, the precision calculation includes all SySLLM key points, encompassing both the points classified in this phase and those already evaluated during the recall assessment. The precision scores, as shown in Table 3, ranged from 0.769 (Dynamic Obstacles agent) to 0.864 (Short-Sighted agent), with an overall mean precision score of 0.819 across all agents. In addition to quantifying correctness, this metric offers a way to evaluate potential hallucinations by identifying points that may not align with expert observations. The high precision scores suggest that hallucinations were minimal in our evaluations. While instances of hallucinated content in SySLLM summaries were rare, they underscore the inherent challenges in summarizing RL agent behaviors using LLMs. One example from the Crossing agent summary: “The agent frequently checks for walls in its path and adjacent tiles.”. This point is classified as “not matched” by all experts, indicating a possible hallucination.
Taken together, the integration of recall and precision metrics demonstrates SySLLM’s capability to produce comprehensive and accurate summaries that closely align with expert observations.
Agent Expert Recall Precision Mean Recall Mean Precision
Unlock Goal-Directed 1 0.500 0.864 0.687 0.864
2 0.643 0.864
3 0.917 0.864
Unlock Short-Sighted 1 0.800 0.846 0.878 0.839
2 0.833 0.807
3 1.000 0.923
Dynamic Obstacles 4 0.583 0.692 0.739 0.769
5 0.833 0.692
6 0.800 0.923
Lava Gap 4 0.667 0.769 0.794 0.811
5 0.786 0.846
6 0.929 0.818
Red Blue Doors 4 0.857 0.767 0.871 0.834
5 0.857 0.867
6 0.900 0.867
Crossing 4 0.750 0.731 0.914 0.795
5 0.917 0.808
6 1.000 0.846
Table 3: Recall and Precision of Expert Insights for SySLLM.
7 User Study
To evaluate the usefulness of the SySLLM summaries, we conducted a user study comparing SySLLM summaries to HIGHLIGHTS-DIV (hereafter referred to as “HIGHLIGHTS”) summaries, which is a standard video-based benchmark in XRL [3]. For the study, we focused on the three agents from the MiniGrid Unlock environment: goal-directed agent, short-sighted agent, and random agent. These agents were chosen to represent qualitatively distinct behaviors, ensuring a diverse set of policies for participants to evaluate. This variety allowed us to better assess how well the summaries captured different decision-making strategies and provided interpretable insights into the agents’ behavior. For completeness, we provide an intuitive explanation of HIGHLIGHTS in Section A of the Appendix. We assessed participants’ subjective preferences as well as their ability to identify agent behavior based on the summaries.
7.1 Procedure
Participants were first shown a tutorial explaining the rules of the Unlock environment. They then had to pass a quiz ensuring they read and understood the rules. Next, they were asked to complete two different tasks.
Task 1: Summary Preferences
In the first task, participants evaluated and compared summaries of agents based on their policies. Initially, they watched a 120-second video showcasing full episodes to familiarize themselves with the agent’s policy. Following this, participants were presented with the first summary, either from the SySLLM or HIGHLIGHTS condition (see Section 7.2), and rated their agreement on a series of metrics using a 7-point Likert scale. The metrics, adapted from Hoffman’s explanation satisfaction questionnaire [15], included Clarity, Understandability, Completeness, Satisfaction, Usefulness, Accuracy, Improvement, and Preference. These adaptations were tailored to the study’s specific needs and excluded irrelevant items (see Section F for the full scale).
Participants then evaluated the second summary, corresponding to the alternative condition (HIGHLIGHTS or SySLLM), using the same set of questions. This allowed for a direct comparison between the two summary types. After assessing both summaries, participants answered additional questions that compared the summaries and expressed their preferences, highlighting which summary they found more helpful and explaining their reasons.
Task 2: Agent Identification
While the first task evaluated participants’ subjective perceptions of the summaries, the second task assessed their objective ability to match a given summary to the correct agent. Participants were presented with summaries, each linked to either the SySLLM or HIGHLIGHTS condition, and shown three short experience videos (approximately 20 seconds each) depicting different agents’ actions in the environment. They were asked to identify which video corresponded to the behavior described in the summary. Participants indicated their chosen video and rated their confidence on a 1-7 Likert scale, along with justifications for their decision in an open text format.
To keep the experiment concise and focused on key comparisons, we set up three matching questions—labeled $Q1$, $Q2$, and $Q3$. Each question featured a summary of a different agent: the goal-directed agent in $Q1$, the random agent in $Q2$, and the short-sighted agent in $Q3$. Depending on the condition to which a participant was assigned, they encountered either a text-based summary (SySLLM) or a video summary (HIGHLIGHTS) for these agents. In each question, alongside the video of the associated agent, participants were also shown two additional videos featuring different, unrelated agents. The task for the participants was to correctly identify the agent that matched the given summary. Success was measured by the participant’s ability to accurately associate the correct agent with its respective summary.
7.2 Experimental Conditions
Participants were divided into four conditions, varying by summary type (SySLLM or HIGHLIGHTS) and agent type (goal-directed or short-sighted). For Task 1, we used a within-subject design where participants viewed summaries of an agent’s behavior in two sequences. To counter ordering effects, one group received the SySLLM summary first and the HIGHLIGHTS summary second, while the other group received the reverse order. In Task 2, we employed a between-subject design. Participants viewed only SySLLM or HIGHLIGHTS summaries, corresponding to the sequence they were exposed to in Task 1. This setup ensured that participants experienced one of four unique combinations of summary and agent types in a controlled manner. Details of these conditions are shown in Table 4.
Benchmark
For generating the HIGHLIGHTS summaries, we used the following parameters: 300 traces (episodes), a context length of 5, and a total of 20 highlights to effectively capture key behaviors and the necessary context within the Unlock environment. These choices ensure that the summaries provide a detailed yet concise depiction of critical actions, appropriately matching the environment’s scale and complexity.
Condition Task 1 Sequence Task 1 Agent Type Task 2 Summary Type
1 SySLLM $\rightarrow$ HIGHLIGHTS Goal-directed Agent SySLLM
2 SySLLM $\rightarrow$ HIGHLIGHTS Short-sighted Agent SySLLM
3 HIGHLIGHTS $\rightarrow$ SySLLM Goal-directed Agent HIGHLIGHTS
4 HIGHLIGHTS $\rightarrow$ SySLLM Short-sighted Agent HIGHLIGHTS
Table 4: Experimental conditions.
7.3 Participants
We recruited participants through Prolific (N = 200). Participants were native English speakers from the US, UK, Australia, or Canada. Participants received a 3.75 pound base payment and an additional, 1 pound bonus if participants answered all questions correctly in Task 2. Participants who failed the attention question were excluded from the final analysis, as well as participants who completed the survey in less time than the length of the videos presented. After exclusions, we had 192 participants (94 female, Mean age = 36.41, STD = 12.05).
7.4 Results
Task 1: Summary Preferences
The SySLLM summaries consistently received higher ratings compared to the HIGHLIGHTS summaries across all metrics (see Figure 3). To assess the statistical significance of the differences in ratings, we first calculate the average scores for each participant under both the SySLLM and HIGHLIGHTS conditions. We then compute the difference between these averages for each participant. A paired t-test is conducted on these differences to evaluate whether the SySLLM condition significantly outperforms the HIGHLIGHTS condition across all metrics. The results indicate a statistically significant difference in favor of the SySLLM condition, with statistic $T=13.99$ and $p<10^{-33}$.
Following the evaluations of specific summaries, we asked the participants to subjectively compare the textual and video summaries. Specifically, we asked them which summary they think is better for understanding the agent’s overall behavior based on the observed video of a large sample of trajectories. They were further asked to rate the extent to which they preferred the textual (or visual) summary (on a 1-7 Likert scale). The results we obtained are conclusive. The responses indicate a clear preference for the SySLLM summary, with 75.5% of participants favoring the language-based summaries. Regarding the second question, the mean rating was 5.97 with a standard deviation of 1.44. In the 1-7 Likert scale, a score of 4 represents indifference to the different policies; hence, a mean score of 5.97 indicates a substantial preference for the textual summary.
Qualitative Analysis The open text feedback we received from participants of Task 1 turned out to be highly valuable. Some participants explained that while HIGHLIGHTS summaries required their own interpretation of the videos, SySLLM summaries provided an interpretation of the agent’s policy, e.g., “the [SySLLM] summary was thorough in explaining the movement of the agent. It describes information that isn’t visible in the video. The information in the summary was objective while watching the video summary and interpreting it is based on my view which is subjective.” Other participants mentioned that SySLLM summaries also explained why the agent acted as it did. For instance, one participant noted that “there are instances in the video where the agent seems to turn random corners instead of straight as you would expect it to. The summary explains why.” Another participant wrote that “the video itself does not fully explain but rather how you interpret what it is doing and in the case of errors it starts doing something wrong you don’t know what causes it”.
Additionally, some participants liked that SySLLM summaries provided more context, e.g., “the textual summary provides context for what is occurring. Conversely, the video summary doesn’t inform the viewer of the ’behind the scenes’ that is occurring. From the textual information, I identified many strategies and code that the agent is using such as using the maze strategy to always turn right when it hits a wall. On the other hand, the video does not provide this information unless you have prior knowledge such as coding experience on conditionals and iterations to know what it occurring.”
Figure 3: Participant scores for Task 1. Each metric includes a box plot of the Likert scores. Notably, SySLLM receives higher ratings than HIGHLIGHTS across all metrics.
Task 2: Agent Identification
As shown in Figure 4, participants’ ability to correctly identify the agent based on the provided summaries varied between the SySLLM and HIGHLIGHTS conditions across the three questions.
For $Q1$ ($\chi^{2}=1.0523$, $p=0.3050$, odds ratio (OR) = 0.5843), $Q2$ ($\chi^{2}=0.6703$, $p=0.4129$, OR = 0.7550), and $Q3$ ($\chi^{2}=0.7279$, $p=0.3936$, OR = 1.3875), the Chi-Square tests showed no statistically significant differences. Independent t-tests showed no statistically significant differences in confidence scores between the SySLLM and HIGHLIGHTS conditions for $Q1$ ($t=-0.6202$, $p=0.5359$) and $Q2$ ($t=0.5565$, $p=0.5785$). However, for $Q3$, there was a statistically significant difference ($t=3.4197$, $p=0.0008$).
To conclude, participants did not show a significant preference for the SySLLM or HIGHLIGHTS condition when identifying the correct agent based on the summaries provided, except for $Q3$ where the SySLLM condition had significantly higher confidence scores. Further investigation may be required to understand the nuances behind these results and whether different summary methods or conditions might yield different outcomes.
Figure 4: Correctness rate in Task 2. For each question, we measure the proportion of users that identified the correct agent under both SySLLM, HIGHLIGHTS, and a random guess. For each bar, we included the 95% confidence intervals. Notice that both SySLLM and HIGHLIGHTS receive higher scores than the random guess. For all three questions, the difference in correctness rate between SySLLM and HIGHLIGHTS is not statistically significant.
8 Discussion and Future Work
We introduce SySLLM, leveraging LLMs to generate summaries of reinforcement learning agents. Its effectiveness is demonstrated across five distinct environments, including seven different agents. Expert evaluations show that SySLLM closely aligns with expert-generated summaries, capturing key aspects of agent behavior. User studies reveal that participants rate SySLLM summaries higher than HIGHLIGHTS on subjective metrics and perform equally well in agent identification tasks. These results provide evidence for the potential usefulness of the proposed approach. We now turn to discuss several limitations of the SySLLM method and possible ways to address them, as well as additional directions for future work.
SySLLM relies on the availability of a captioning function, a widely accepted approach in language-integrated methodologies, as evidenced by prior research [13, 26, 20]. In simulated environments, captioning is particularly straightforward, as ground truth states are readily accessible, and numerous RL environments, such as NetHack/MiniHack [32, 22] and text-based games [35, 11], already incorporate language-based features. In more complex or real-world scenarios, robust techniques such as object detection [41], cutting-edge captioning frameworks [36], action recognition systems [21], or advanced vision-language models [2] can be utilized. Alternatively, annotator models [13] or pretrained foundation models [37] offer viable solutions for generating captions. With continued advancements in general-purpose captioning models, we anticipate that readily available, off-the-shelf captioning tools will soon support an even broader range of tasks, further solidifying the practicality of this approach.
Advancements in LLMs, particularly larger context windows, can extend this approach to large-scale and more complex environments, further validating its effectiveness and unlocking its potential to provide deeper insights into agent behavior.
Future work could explore a question-answering (QA) framework, allowing users to query the agent’s policy directly. For example, users could ask why specific actions were taken or predict the agent’s behavior in new scenarios. This interactive system would provide deeper insights into the agent’s decision-making process, offering a flexible and intuitive way to understand its behavior.
Another avenue for future work is generating multimodal summaries, combining natural language explanations with visual representations of key moments, trajectories, and decisions. That is, for each pattern identified by the LLM, the summary would also include demonstrations of the pattern extracted from agent traces. This approach could enhance user understanding by integrating text and visuals to more effectively convey complex policies. Visual elements from the agent’s experience dataset might help mitigate hallucinations by grounding explanations in observable behavior, potentially improving reliability and trustworthiness.
Acknowledgments
This project was funded by the ERC Starting Grant #101078158 CONVEY and the Israel Science Foundation (ISF) under Grant No. 3079/24.
Appendix A HIGHLIGHTS
“Highlights” Policy Summaries
Our user study uses “Highlights” policy summaries [3] as a baseline. For completeness, we describe the algorithm here. The HIGHLIGHTS algorithm generates an online summary of an agent’s behavior from simulations, using state importance to decide which states to include. A state is considered important if taking a wrong action there significantly decreases future rewards, as determined by the agent’s Q-values. Formally, state importance $I(s)$ is defined as:
$I(s)=\max\limits_{a}Q^{\pi}_{(s,a)}-\min\limits_{a}Q^{\pi}_{(s,a)}.$
HIGHLIGHTS captures trajectories with the most important states encountered in simulations. At each step, it evaluates state importance and adds the state to the summary if its importance exceeds the current minimum in the summary, replacing the least important state. For each state added, it also includes a trajectory of neighboring states and actions.
To address redundancy in similar important scenarios, the HIGHLIGHTS-DIV algorithm extends HIGHLIGHTS by incorporating diversity. HIGHLIGHTS-DIV evaluates a state $s$ by identifying the most similar state $s^{\prime}$ in the summary. It compares $I(s)$ to $I(s^{\prime})$ instead of the minimum importance value. If $I(s)$ is greater, the trajectory including $s^{\prime}$ is replaced with the current trajectory. This approach maintains less important but diverse states, enhancing the information conveyed to users.
Appendix B Information Stored in Experience dataset
Information Description
Episode Number The number of the episode from which the data was collected.
Step Number The specific step within the episode.
Captioned Observation The observation converted into natural language.
Captioned Action The action converted into natural language.
Cumulative Reward The total reward accumulated by the agent up to that step.
Table 5: Description of the data stored in the experience dataset.
Appendix C Summarization Prompt:
[General Instructions]
Generate a focused summary of the RL agent’s policy based on the provided episodes data. Highlight key behaviors, decision-making processes, and patterns specific to this agent. Tailor the summary to reflect unique strategies and actions observed.
Focus on:
• Recurring patterns and behaviors specific to this agent’s policy.
• Detailed analysis of decision-making processes and responses to different stimuli.
• Efficiency in identifying and interacting with relevant objects (e.g., keys, doors).
• Methods used to solve tasks and handle obstacles.
• Comparison of agent’s performance across different episodes.
• Quantitative metrics (e.g., number of steps, success rates) to evaluate efficiency.
• Analysis of navigation strategies and adaptations to the environment.
• Provide insights on the variability and randomness of the agent’s actions and decisions. Look at the distribution of the agent’s actions during the episodes.
• Describe any notable fluctuations or inconsistencies in behavior across episodes.
• Figure out what is the agent’s field of vision and its influence on its decisions and actions, considering it may not see the entire grid but only a part of it.
[Environment Description]
• Goal: Maximize cumulative reward (close to 1) by opening the door.
• The environment is static, with only the agent moving within it. Walls are fixed and form the boundaries of the grid (there are no further obstacles).
• Possible agent’s actions: move forward, turn left, turn right, pick up, toggle, drop.
• Max episode length: 70 steps (failure to open the door if exceeded).
[Interpretation Instructions]
• If the next action is blank, the episode ended successfully. Use the cumulative reward to analyze performance.
• Follow the agent’s received observations to understand its behavior and decision-making process.
• Performance metrics: Number of steps to open the door, cumulative reward.
• Descriptions are from the perspective of the agent, with the number of tiles indicating the distance.
[Output Instructions]
Provide a concise summary focused on specific behaviors and actions of the agent. Formatted for easy understanding by non-technical individuals. Avoid general or trivial statements; focus on specific insights from the agent’s behavior. Include quantitative metrics and comparisons to provide a clear understanding of performance.
[Formatted Experience Dataset]
Appendix D Unlock Goal-Directed Agent Full Summary
This is an example of a summary produced by the SySLLM algorithm, illustrated here using the Unlock goal-directed agent:
The analysis of the RL agent’s behavior across 70 episodes reveals a consistent pattern in its decision-making and interaction with the environment. The agent’s primary goal is to locate and use a key to open a corresponding colored door within a grid environment. Here are the key observations and insights from the agent’s performance:
1. Navigation and Object Interaction: - The agent effectively identifies and navigates towards keys, demonstrating a capability to adjust its path based on the key’s relative position. Once the key is picked up, the agent shifts its focus to finding and reaching the corresponding door. - The agent uses a combination of ’move forward’, ’turn left’, and ’turn right’ actions to navigate towards objectives, showing proficiency in altering its path dynamically based on the environment’s layout and the objects’ positions.
2. Efficiency Metrics: - The agent’s efficiency varies across episodes, with some episodes achieving the goal in fewer steps and higher cumulative rewards. - The average steps taken to complete an episode when successful are approximately 15-25 steps, with cumulative rewards often nearing the maximum, indicating high efficiency in several scenarios.
3. Decision-Making Patterns: - The agent demonstrates a pattern of turning towards the nearest key or door once identified, suggesting a straightforward heuristic of minimizing distance to the target. This behavior is consistent across episodes, indicating a robust underlying policy for basic navigation and objective fulfillment. - In scenarios where the agent is positioned with immediate obstacles or walls, it shows a tendency to turn and reorient itself towards open paths, reflecting an ability to avoid simple traps and dead ends.
4. Handling of Obstacles: - The agent effectively navigates around walls and adjusts its path to reach its targets. There is no evidence of the agent being stuck or failing to navigate around corners, which suggests good spatial awareness and pathfinding within the grid constraints.
5. Variability and Randomness: - The agent’s actions show some variability, particularly in scenarios where multiple paths can lead to the objective. This variability might indicate a level of exploration or randomness in the policy, allowing the agent to discover and adapt to new optimal paths.
6. Performance Consistency: - Across different episodes, the agent’s performance shows consistency in terms of strategy but varies in execution efficiency, likely due to the differing initial conditions and environmental layouts. The agent consistently attempts to pick up the nearest key and unlock the door, aligning with its primary objective.
7. Field of Vision Impact: - The agent’s decisions are heavily influenced by its immediate field of vision, as it tends to react to objects as they become visible. This suggests a limited perception range, requiring the agent to physically move closer to distant objects to detect and interact with them.
In summary, the RL agent is proficient in navigating grid environments to achieve specified goals of picking keys and opening doors. It demonstrates effective use of basic navigational actions, adapts well to the visible environment, and shows a consistent approach to achieving its objectives, albeit with some variability in efficiency possibly due to explorative behaviors or initial positioning. The agent’s performance could potentially be enhanced by expanding its field of vision or optimizing its pathfinding algorithms to increase consistency in achieving higher rewards in fewer steps.
Appendix E Implementation Details
We employed the PPO algorithm from the stable-baselines3 library for our policy network, which takes as input a $K\times K\times 3$ encoded image and a mission string, the latter being encoded using a one-hot scheme. These inputs are combined into a single 2835-dimensional vector. The network architecture features two hidden layers, each comprising 64 neurons, with ReLU activation functions introducing non-linearity. The output layer, designed to match the 6-dimensional action space of the environment, utilizes a softmax activation function to generate a probability distribution over possible actions. Additionally, we normalized the observations. For the short-sighted agent, the observation grid size is $3\times 3\times 3$, while for the goal-directed agent, it is $11\times 11\times 3$.
Hyperparameter Goal-Directed Short-Sighted Dynamic Obstacles Lava Gap Red Blue Doors Crossing
Total Timesteps $2\times 10^{6}$ $1\times 10^{6}$ $2\times 10^{6}$ $2\times 10^{6}$ $2\times 10^{6}$ $3\times 10^{6}$
Number of Environments 8 8 8 16 8 16
Number of Steps 512 512 2048 1024 512 2048
Batch Size 64 64 256 128 64 256
GAE Lambda (gae_lambda) 0.95 0.95 0.95 0.95 0.95 0.95
Discount Factor (gamma) 0.99 0.99 0.99 0.99 0.99 0.99
Number of Epochs 10 10 30 10 10 20
Entropy Coefficient 0.001 0.001 0.01 0.001 0.001 0.01
Learning Rate $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
Clip Range 0.2 0.2 0.2 0.2 0.2 0.2
Table 6: Hyper-parameters for the PPO algorithm applied to all six agents.
Appendix F Experts instructions:
General Instructions:
Generate a focused summary of the RL agent’s policy based on the provided episodes data. Highlight key behaviors, decision-making processes, and patterns specific to this agent. Tailor the summary to reflect unique strategies and actions observed.
Focus on:
  • Recurring patterns and behaviors specific to this agent’s policy.
  • Detailed analysis of decision-making processes and responses to different stimuli.
  • Efficiency in identifying and interacting with relevant objects (e.g., keys, doors).
  • Methods used to solve tasks and handle obstacles.
  • Comparison of agent’s performance across different episodes.
  • Quantitative metrics (e.g., number of steps, success rates) to evaluate efficiency.
  • Analysis of navigation strategies and adaptations to the environment.
  • Provide insights on the variability and randomness of the agent’s actions and decisions. Look at the distribution of the agent’s actions during the episodes.
  • Describe any notable fluctuations or inconsistencies in behavior across episodes.
  • Figure out what is the agent’s field of vision and its influence on its decisions and actions, considering it may not see the entire grid but only a part of it.
Environment Description:
  • Goal: Maximize cumulative reward (close to 1) by opening the door.
  • The environment is static, with only the agent moving within it. Walls are fixed and form the boundaries of the grid (there are no further obstacles).
  • Possible agent’s actions: move forward, turn left, turn right, pick up, toggle, drop.
  • Max episode length: 70 steps (failure to open the door if exceeded).
Summary Instructions:
The agent description should be at least 100 words. Provide approximately 5 key insights.
Appendix G Scale Used in Task 1
In Task 1 of our study, we utilized a 7-point Likert scale to evaluate participants’ perceptions and understanding of the agent’s behavior as presented in both the video summaries and the natural language summaries. Participants rated their agreement with the following statements, where 1 indicates “Strongly disagree” and 7 indicates “Strongly agree”. The questions were phrased according to the condition—either video or natural language summary.
  1. 1.
    Clarity: “The [video/natural language] summary clearly explained the agent’s actions and decisions shown in the demonstration video.”
  2. 2.
    Understandable: “From the [video/natural language] summary, I understand how the agent’s actions and decisions shown in the demonstration video.”
  3. 3.
    Completeness: “The [video/natural language] summary seemed complete in covering all aspects of the agent’s actions and decisions in the demonstration video.”
  4. 4.
    Satisfaction: “The [video/natural language] summary is satisfying in capturing the agent’s behavior and decisions displayed in the demonstration video.”
  5. 5.
    Useful: “The [video/natural language] summary is useful to my understanding of the agent’s behavior and decisions displayed in the demonstration video.”
  6. 6.
    Accuracy: “The information in the [video/natural language] summary accurately reflected the agent’s behavior and decisions displayed in the demonstration video.”
  7. 7.
    Improvement: “The [video/natural language] summary provides additional insights about the agent’s behavior that are not immediately apparent from watching the demonstration video alone.”
  8. 8.
    Preference: “I prefer receiving information about agent behavior through the [video/natural language] summary rather than just watching the demonstration video.”
These ratings provided quantitative data to assess the effectiveness and clarity of both the video and natural language summaries in conveying the agent’s behavior and decision-making processes. This scale aimed to capture various dimensions of participant satisfaction and understanding, contributing to the overall evaluation of the summaries’ utility in the context of our research.
Appendix H Systematic Exploration of the Prompt Design
The creation of the final prompt was achieved through a structured and iterative exploration process. This process involved a quantitative evaluation of prompt designs based on observed outputs, guided by principles from prompt engineering literature, and tailored to domain-specific requirements. Additionally, the final design was inspired by the Chain of Thought (CoT) [40] prompting paradigm, which encourages models to generate structured, step-by-step reasoning. Below is a detailed breakdown of the methodology used:
Define the Objective
Goal: The primary objective of the prompt was to generate a focused and comprehensive global summary of the policy of the RL agent. The summary needed to highlight key behaviors, decision-making processes, and performance metrics in a manner understandable to both technical and non-technical audiences, while ensuring it could function as a zero-shot prompt without requiring additional training examples.
Key Constraints:
  • The prompt must guide the model to produce specific, concise, and informative summaries.
  • It should minimize general or trivial statements and focus on insights from the agent’s behavior.
Decomposition of Requirements
To meet the objective, the task was broken down into several core components:
  • Behavioral Analysis: Capturing recurring patterns, strategies, and responses to stimuli.
  • Performance Metrics: Including quantitative insights such as success rates and steps taken.
  • Environmental Factors: Reflecting the influence of the agent’s field of vision and static surroundings.
  • Comparison Across Episodes: Addressing variability and randomness in actions.
  • Accessibility: Ensuring the output is clear and digestible for non-technical readers.
Iterative Prompt Design
Initial Prototype:
  • Focused on general instructions for summarization.
  • Included high-level tasks such as “describe the agent’s behavior” without specifying details.
Issues Identified:
  • Outputs were overly generic, lacked depth, and failed to focus on specific behaviors or metrics.
Refinement 1: Add Specific Focus Areas
  • Incorporated bullet points to guide the model to focus on particular aspects, such as “recurring patterns,” “quantitative metrics,” and “navigation strategies.”
Observations:
  • Improved relevance and depth of the summaries.
  • However, the outputs lacked consistency in formatting and interpretability.
Refinement 2: Structured Prompt Sections
  • Segmented the prompt into distinct parts:
    • General Instructions
    • Environment Description
    • Interpretation Instructions
    • Output Instructions
    • Formatted Experience Dataset
Observations:
  • Enhanced structure improved consistency.
  • More detailed context in “Environment Description” provided clarity for the model to ground its responses.
Refinement 3: Inspired by Chain of Thought (CoT) Reasoning
  • The prompt was designed to encourage a step-by-step analysis, mirroring the CoT paradigm:
    • Each bullet point and section was treated as a sub-task requiring focused attention.
    • For example, instructions like “Analyze navigation strategies and adaptations to the environment” explicitly directed the model to break down its reasoning into smaller, manageable steps.
Observations:
  • Outputs exhibited improved logical flow and comprehensive coverage of required aspects.
  • The structured approach mitigated issues with overly generic or shallow responses.
Refinement 4: Emphasize Quantitative and Comparative Analysis
  • Added explicit instructions to include metrics like “number of steps” and “success rates.”
  • Introduced the requirement to compare the agent’s performance across episodes.
Observations:
  • Summaries became more data-driven and analytical.
  • Increased attention to variations in the agent’s behavior.
Refinement 5: Addressing Accessibility
  • Adjusted language in the “Output Instructions” to ensure summaries were understandable to non-technical audiences.
  • Included a directive to avoid trivial statements.
Final Testing:
  • Conducted multiple test runs with varied episode datasets.
  • Evaluated the prompt’s ability to guide the model toward producing outputs that met the objective.
  • Fine-tuned phrasing for clarity and focus.
Key Design Considerations
Clarity and Specificity:
  • Each section of the prompt was crafted to minimize ambiguity, ensuring the model understood the task requirements.
Structure Inspired by CoT:
  • The step-by-step breakdown mirrored the CoT prompting approach, which is known to improve reasoning and response quality in large language models.
Focus on Insightful Analysis:
  • By explicitly asking for “variability,” “distribution of actions,” and “quantitative comparisons,” the prompt steered the model toward generating meaningful insights.
Evaluation and Lessons Learned
Evaluation:
  • Outputs were analyzed for relevance, specificity, and clarity.
  • Feedback from test runs informed iterative improvements.
Lessons Learned:
  • Prompts benefit from structured sections that provide clear and detailed guidance.
  • Incorporating CoT-inspired design principles encourages logical, step-by-step reasoning in outputs.
  • Tailoring language for accessibility improves utility for non-technical audiences.
Rationale for the Final Design
The final prompt integrates the following elements:
  • Comprehensive Instructions: Ensuring detailed and targeted outputs.
  • Quantitative Focus: Providing measurable insights for evaluating agent performance.
  • Clarity and Accessibility: Catering to a broad audience, including non-technical users.
  • Structure Inspired by CoT: Encouraging the model to follow a logical sequence in generating summaries.
This systematic process, incorporating insights from the Chain of Thought paradigm, demonstrates the thoughtful process taken to ensure the prompt is both effective and robust for summarizing RL agent policies.